Find the present values of these ordinary annuities. Discounting occurs once a year. Do not round intermediate calculations. Round your answers to the nearest cent.

$700 per year for 12 years at 14%.

$350 per year for 6 years at 7%.

$400 per year for 6 years at 0%.

Rework previous parts assuming they are annuities due. Present value of

$700 per year for 12 years at 14%:

$ Present value of $350 per year for 6 years at 7%:

$ Present value of $400 per year for 6 years at 0%:
Answer #1

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

1.Present value=700[1-(1.14)^-12]/0.14

=700*5.660292125

=$3962.20(Approx).

2.Present value=350[1-(1.07)^-6]/0.07

=350*4.76653966

=$1668.29(Approx).

3.Present value=400*6

=$2400

Present value of annuity due=Present value of annuity*(1+rate)

1.Present value=3962.20*1.14

=$4516.91(Approx).

2.Present value=1668.29*1.07

=$1785.07(Approx).

3.Present value=400*6

=$2400